OpenStudy (anthonylocke):

I'm doing roots and radicals. 2√12 = 4√3 Are these equals?

OpenStudy (anthonylocke):



OpenStudy (anthonylocke):

That's the problem.

OpenStudy (little_miss):

yes they are equal

OpenStudy (sshayer):

yes \[2\sqrt{12}=2\sqrt{4 \times3}=2 \times2\sqrt{3}=4\sqrt{3}\]

OpenStudy (anthonylocke):

thanks, i really appreciate it.

OpenStudy (sshayer):

yw

OpenStudy (anthonylocke):

do i give you guys a point or something? I'm not familiar with this site.

OpenStudy (sshayer):

just click best response

OpenStudy (little_miss):

yeah you can give us both a best response